Diffusion Morning Report 03/10/11

Ted Berg tberg at stanford.edu
Thu Mar 10 09:56:15 PST 2011


Good Morning All,
     Looking good this morning. Epi is back together and being checked 
out. Thermco LTO is back and will be installed. TEOS pump and Cryo 
compressor out for repair. Tylan 1 power fail on MFS 460. Thermco 2 no 
tlc flow. No other major problems at this time.

					
	Date: 	3/10/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*down* 	*power faIL* 	** 	** 	*2*
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*pump bad-out for repair* 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*20.35/30* 	** 	**
Tylan 10-Nitride 	*yellow* 	*door close issue* 	*7.89/30* 	** 	**
Tylan 11-LTO 	*up* 	*a bit thin* 	*21.71/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*down* 	*bad pump-out for repair* 	** 	** 	*3*
AMAT Epi 	*down* 	*re-assembled need conditioning and CO* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor out for repair* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 673 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110310/6629fc48/attachment.html>


More information about the diffusion mailing list